### v2

# Scaling laws for post-training quantized large language models

Abstract
Generalization abilities of well-trained large language models (LLMs) are known to scale predictably as a function of model size. In contrast to the existence of practical scaling laws governing pre-training, the quality of LLMs after post-training compression remains highly unpredictable, often requiring case-by-case validation in practice. In this work, we attempted to close this gap for post-training weight quantization of LLMs by conducting a systematic empirical study on multiple LLM families quantized to numerous low-precision tensor data types using popular weight quantization techniques. We identified key scaling factors pertaining to characteristics of the local loss landscape, based on which the performance of quantized LLMs can be reasonably well predicted by a statistical model.
1 Introduction
Large language models (LLMs) based on the transformer architecture (Vaswani et al., 2023) are known to obey empirical scaling laws. An LLM’s generalization abilities, measured by the negative-log-likelihood (NLL) loss in next-token prediction, are predictably related to increases in parameter count, pre-training data volume, and computation cost (Kaplan et al., 2020; Dettmers and Zettlemoyer, 2023; Henighan et al., 2020; Alabdulmohsin et al., 2022; Su et al., 2024; Song et al., 2024; Muennighoff et al., 2023; Bordelon et al., 2024; Bahri et al., 2024).
Thanks to the guidance from these scaling laws, pre-training of LLMs, a notoriously expensive computation in practice, enjoys a certain degree of confidence in return on investment. However, for these LLMs to run efficiently on a target accelerator for inference, they often need to undergo post-training compression, such as quantization (Gholami et al., 2021; Frantar et al., 2022; Park et al., 2024; Kim et al., 2023, 2024; Yao et al., 2022).
Post-training quantization (PTQ) is a process that attempts to preserve a trained LLM’s generalizability, while performing its computation with low-precision data types. Because PTQ involves many additional factors, it introduces significant uncertainty into the quality of the final model, in many cases completely obscuring the predictability prescribed by the pre-training scaling laws. This makes PTQ a business of trial-and-error (Huang et al., 2024; Sharify et al., 2024; Yuan et al., 2023; Hu et al., 2022), lacking the useful practical guidance from scaling laws like those that govern pre-training.
Beyond just algorithmic complexity, PTQ also becomes incredibly time and compute intensive when one attempts to find the optimal quantization format and model parameter count given fixed memory, compute, and data format constraints. A simple illustration of this trade-off is the comparison between a larger model quantized to a lower bit format and a smaller model quantized to a higher bit format, a search space that requires a lot of iterations, and consequently, significant time and compute.
In this work, we attempted to close this gap in knowledge by systematically studying the empirical scaling of extra factors involved in PTQ in addition to the pre-trained NLL loss. We briefly enumerate below all factors considered.
  1. 1.
    Loss of pre-trained LLM. A known scaling law governs the relationship between LLM training parameters and the quality of the resulting model. Intuitively, a better trained model would also have better performance in a quantized state, so the initial loss of a pre-trained LLM is highly relevant to profiling the quantized loss landscape. Section 2.1 is dedicated it.
  2. 2.
    Local loss landscape of pre-trained LLM. Because quantization is a specific perturbation to the trained network, the resulting loss due to the perturbation depends not only on the converged NLL loss, but also on how steeply the loss changes in the neighborhood of convergence  (Frumkin et al., 2023; Nahshan et al., 2020; Evci et al., 2020). Section 2.2 is dedicated to understanding how the local loss landscape changes with scale.
  3. 3.
    Low-precision data type for quantization. Numerous novel tensor data types for efficient inference have emerged recently (Rouhani et al., 2023; Dettmers et al., 2023; Agrawal et al., 2024; Guo et al., 2022). Intuitively, both the tensor data type and its numerical precision would correlate with the quality of quantization, and Section 2.3 is dedicated to its scaling.
  4. 4.
    PTQ algorithm. After aggressive low-precision quantization, certain PTQ optimization algorithms are commonly used to recover some model quality (Frantar et al., 2022; Xiao et al., 2024; Lin et al., 2024; Lee et al., 2024). These methods typically minimizes local quantization error as opposed to direct global loss optimization as in quantization-aware fine-tuning (e.g. Li et al. 2023; Jeon et al. 2024). Section 2.4 is dedicated to profiling how those properties scale.
We show with concrete examples (for procedural details see Section 4), that all the above factors have underlying empirical scaling laws for certain LLM families. Incorporating these empirical rules, in Section 3, we build a predictive statistical model that takes the above factors as input and predicts the outcome of a PTQ procedure on unseen LLMs at a reasonable accuracy.
2 Factors subject to scaling for LLM PTQ
Figure 1: Left: Scaling of pre-trained NLL loss. NLL losses evaluated on the validation split of the WikiText-2 dataset are plotted against the total parameter counts in the transformer layers’ weight tensors. Model families are color-coded and the symbol sizes encode the weight parameter count, a convention shared by following figures. Right: Local radial loss landscape mapping. Shown here is measurement of the typical loss landscape in the neighborhood of pre-trained weights, by evaluation of the loss along typical radial perturbations, 3 independent instances illustrated for opt-1.3b, together with their Taylor series approximations.
Figure 2: Left: Local loss landscape of LLMs grouped in families. Shown are the mean (colored curves) and range (colored shades) of 3 independent measurements for each model. The typical characteristics are common to all models. Within a family, larger models tend to have flatter local loss landscape, in a predictable manner. Right: Scaling of local loss landscape as a function of LLM size. We plot NLL loss against weight parameter count, with typical perturbation SNR as a gray-scale heat map. Thin white iso-SNR curves are at 2 dB increments. With OPT family as the only exception, vertical spacing of these iso-SNR curves is shorter in large models than in small ones of the same family, suggesting flatter local minima at larger model sizes.
2.1 Loss of pre-trained LLM
First, we recapitulate one of the original scaling laws on well trained LLMs with no data limit (Kaplan et al., 2020). We visualize in Figure 1 (left) this scaling law with our experiments (see Section 4 for details). The GPT-2, OPT and BLOOM model families roughly follow one power law, whereas models in the Llama 2/3 family track a different, but qualitatively similar path.
2.2 Characteristics of local loss landscape
Next, we characterize another crucial factor intrinsic to the LLM itself, its local loss landscape.
A quantization of network weight ${\bm{w}}$  [^1] Here we denote by vector ${\bm{w}}$ a flattened version of all weight matrices $\left({\bm{W}}_{1},\cdots,{\bm{W}}_{L}\right)$ of the network that are subject to quantization. can be considered as a perturbation ${\bm{w}}\rightarrow{\bm{w}}+\Delta{\bm{w}}=Q({\bm{w}})$, where $Q$ is a quantizer, and the resulting loss of the quantized network becomes $\mathrm{NLL}({\bm{w}}+\Delta{\bm{w}})$ from the pre-trained $\mathrm{NLL}({\bm{w}})$. The resulting loss is a function not only of the pre-trained weight ${\bm{w}}$, but also of the perturbation $\Delta{\bm{w}}$, often approximated by Taylor expansion,
$\mathrm{NLL}({\bm{w}}+\Delta{\bm{w}})$ $=\mathrm{NLL}({\bm{w}})$
$+{\bm{g}}^{\top}\Delta{\bm{w}}+\frac{1}{2}\Delta{\bm{w}}^{\top}{\bm{H}}\Delta{\bm{w}}$
$+O(\left\lVert\Delta{\bm{w}}\right\rVert^{2}).$
[…]
As the absolute magnitude of ${\bm{w}}$ scales with dimensionality (see Appendix A), we use signal-to-noise ratio (SNR), a relative quantity to measure the magnitude of its perturbation $\Delta{\bm{w}}$,
$\mathrm{SNR}({\bm{w}},\Delta{\bm{w}})$ $=20\log_{10}\frac{\left\lVert{\bm{w}}\right\rVert}{\left\lVert\Delta{\bm{w}}\right\rVert},$
in decibel (dB). A higher SNR represents a smaller deviation $\Delta{\bm{w}}$ from ${\bm{w}}$. When the perturbation is due to quantization, i.e. $\Delta{\bm{w}}=Q({\bm{w}})-{\bm{w}}$, SNR becomes signal-to-quantization-noise ratio (SQNR),
$\mathrm{SQNR}({\bm{w}})$ $=20\log_{10}\frac{\left\lVert{\bm{w}}\right\rVert}{\left\lVert Q({\bm{w}})-{\bm{w}}\right\rVert}.$
Intuitively, the flatter the local loss landscape is near ${\bm{w}}$, the less impact a same perturbation $\Delta{\bm{w}}$ is to exert on the loss. In Figure 1 (right), we show with an example LLM, the typical local loss landscape in the neighborhood of pre-trained weights. We randomly sample a unit vector $\hat{\bm{e}}\sim{\mathcal{S}}^{D}$ from the $D$-dimensional unit sphere, $D$ being the dimensionality of ${\bm{w}}$, and measure $\mathrm{NLL}({\bm{w}}+\lambda\hat{\bm{e}})$ while sweeping $\lambda\in{\mathbb{R}}^{+}$. We see that the typical radial loss is very step-like: it stays relatively low and flat near ${\bm{w}}$, then rises rapidly (faster than quadratic), and finally plateaus further away from ${\bm{w}}$. These qualitative characteristics are shared by all LLMs of various sizes and from various families (Figure 2, left).
We also find that, within the same LLM family, larger models have flatter local loss landscape than smaller ones, in a systematic way (Figures 2) for each family.
2.3 Low-precision data type for quantization
Now, we identify an extrinsic factor in PTQ process: the low-precision tensor data type for quantization. Note that we consider tensorial data types, not simply scalar numerical formats. In addition to traditional integer quantization that requires calibration, emerging standards such as microscaling (MX, Rouhani et al. 2023) adopt more effective and efficient tensor data types, which we study in this work. we also present a comparative study of traditional integer quantization in Appendix C.
We first ask how the magnitude of quantization errors $\Delta{\bm{w}}=Q({\bm{w}})-{\bm{w}}$ vary across LLMs for certain data types. Despite the existence of significant scaling of $\left\lVert{\bm{w}}\right\rVert$ (see Appendix A for further details), the SQNRs are relatively invariant across model families and model sizes, and vary across numerical data types in a highly predictable manner (see details in Appendix  B). In contrast, NLL losses show a much more nonlinear and less predictable pattern, with a rough trend of lower precision data formats leading to higher losses (see details in Appendix  B).
However, with certain choices of weight data type, the perturbation due to quantization is significantly flatter than the typical flatness of the local loss landscape, which we shall elaborate in the next section.
Figure 3: Left: Scaling of SQNRs and NLL losses before and after PTQ, relative to the typical loss landscape. We show data from 3 members of the OPT model family, whose parameter counts are separated by 1 order of magnitude. RTN (before PTQ, hollow symbols) and GPTQ (after PTQ, filled symbols) are plotted together with the typical radial loss landscape empirically mapped. Right: Local loss landscape underlying varied effectiveness of GPTQ acting on the same model quantized at different weight precision. Shown here are data of opt-1.3b quantized to mxint6_128, mxint4_128, mxint3_128 and mxint2_128. The colored, hollow or filled diamonds represent the SQNRs and NLL losses before and after GPTQ, respectively. We further map the underlying radial loss landscape in the directions of typical random perturbation (thin gray lines), of RTN quantization (colored dashed lines) and of GPTQ quantization (colored solid lines).
Figure 4: Changes in SQNRs and NLL losses resulting from GPTQ for OPT family. Numerical precision is color-coded and model size encoded by symbol size. Diagonal line represents identity.
2.4 PTQ optimization method
Finally, we study another important extrinsic factor that contributes to the quality of quantized LLMs for inference, the PTQ optimization algorithm.
To each model and for each weight data type, we applied an improved GPTQ procedure (see Section 4.3 for details) to further optimize the RTN quantized network. Figure 3 (left) shows 3 members of varied sizes from the OPT family. Apparently, the application of GPTQ generally reduced both the SQNR and NLL loss of the RTN model. The reduction in SQNR is relatively consistent across model sizes and data formats, whereas the reduction in NLL loss is highly variable as a function of model size and quantization precision in, however, a rather systematic way. An aggregation of direct comparisons of SQNRs and NLL losses before and after the GPTQ procedure for the OPT model family is presented in Figure 4.
With our systematic collection of empirical data pertaining to all the above-mentioned factors, we are able to uncover patterns in the highly varied, and seemingly haphazard, effect of GPTQ on given a specific LLM quantized to a specific numerical data type. Here we demonstrate with the model opt-1.3b subject to quantization to mxint6_128, mxint4_128, mxint3_128 and mxint2_128 (Figure 3, right). The observation is that GPTQ greatly improves mxint3_128 quantization, but only marginally improves its 6-bit, 4-bit and 2-bit counterparts. The effect of GPTQ seems highly non-monotonic as a function of quantization precision. Nevertheless, in the light of the underlying local loss landscape, the phenomenon can be well understood. First, RTN quantization to MX weight formats often lead to perturbations that are flatter than typical radial loss profiles; the application of GPTQ, further seeks an even flatter perturbation direction in the loss landscape, as evident in Figure 3 (right). However, because these radial loss profiles are very step-like, any linear or quadratic approximations typically fail to characterize them well at SNRs lower than 20 dB. Because of the difference in the effective radii between the RTN and GPTQ loss profiles that are both step-like, a narrow window in SNR exists within which the effect of GPTQ is substantial. Note that the location and size of this window is a function of the model family, the model size, and the numerical data type for weight quantization, as we described above.
Figure 5: Left: A predictive model based on random forest regression. Data for 18 models from the 5 LLM families used for predictive model fitting are shown in light gray; colored symbols represent held-out test data from mpt-7b and pythia-1b, respectively. Prediction and observation are plotted against each other for direct comparison, diagonal line marking identity. Right: Prediction of NLL losses after GPTQ, for unseen LLMs. We tested our predictive model’s performance on 2 held-out LLMs from unseen model families, mpt-7b and pythia-1b. Convention follows Figure 3, with additional large circular symbols representing model prediction of GPTQ losses.
3 Building a predictive model
To sum up our findings thus far, we first found that the characteristics of local loss landscape, just like the loss itself, scales with model size in LLM families, an intrinsic model property. We also determined that choices of the low-precision data type for quantization and the PTQ process, acting within the local loss landscape, lead to different SQNRs and losses in a predictable way.
Taking these empirical rules into consideration, we now build a predictive model based on random forest regression. We set the hyperparameters, the number of estimators and maximum depth of the regressor, to 120 and 8, respectively. The regressor takes a few empirically measured features as input, and directly predicts the resulting NLL loss of the final, quantized model. Given a specific LLM and a specific MX data format with quantizer $Q$, the input features are: (a) weight parameter count $D$, (b) pre-trained loss $\mathrm{NLL}({\bm{w}})$, (c) SQNR of RTN quantization $\mathrm{SQNR}({\bm{w}})$, (d) loss of RTN quantization $\mathrm{NLL}\left(Q({\bm{w}})\right)$, (e) radial slope of local loss landscape at RTN weights $\frac{d\mathrm{NLL}}{d\mathrm{SQNR}}\big{|}_{Q({\bm{w}})}$, (f) numerical format’s precision P, number of element exponent bits E, and block size K. The model outputs a predicted loss after GPTQ, $\mathrm{NLL}(Q({\bm{w}}^{*}))$.
We fit the model on all feature data collected from models in the 5 LLM families above, and test its prediction for 2 held-out models from unseen model families, namely EleutherAI/pythia-1b and mosaicml/mpt-7b. Despite the difference in model architecture, training paradigm, and even local loss landscape between pythia-1b, mpt-7b, and our existing model families, the prediction is reasonably accurate (Figure 5), suggesting that the underlying scaling laws are generalizable across both different model sizes and different LLM families. See Appendix D for detailed interpretation of the predictive model and salient features.
4 Experimental procedures
4.1 Models and dataset
We experimented with models from 5 LLM families, namely GPT-2 (Radford et al., 2019), OPT (Zhang et al., 2022), BLOOM (Workshop et al., 2023), Llama 2 (Touvron et al., 2023), and Llama 3 (Meta, 2024). The models were served by the Hugging Face Model Hub. We identify the models by their unique name string identifier throughout this paper, with their organization prefixes sometimes omitted for brevity.
To validate the generalizability of our empirical scaling rules extracted from studying the above 5 model families, we tested their predictive power on 2 held-out LLMs, EleutherAI/pythia-1b (Biderman et al., 2023), and mosaicml/mpt-7b (MosaicML, 2023).
The WikiText-2 dataset (Merity et al., 2016) was used in all experiments, with the text tokenized by corresponding tokenizers at maximum sequence length of each respective model. 128 examples from the training split were used as calibration dataset for PTQ algorithms. All examples from the validation split were used for validation.
4.2 Numerical tensor data type and notations
We experimented with microscaling (MX, Rouhani et al. 2023) compliant data formats, where a block of tensor elements share a same scaling factor in the format of e8m0 (8-bit exponent and 0-bit mantissa), and each element being of a low-precision float or int number. We experimented with 36 distinct MX data types with precision with block sizes ranging from 16 to 128, and element precision from 2 to 6.
We denote MX formats by mxfpP_eEmM_K or mxintP_K, following the notation from community standard (Rouhani et al., 2023), where P is the precision, K the block size, and E, M the numbers of element exponent and mantissa bits. For example, mxint6_64 represents an MX data type where the element is in int6 and the block size 64; mxfp4_e2m1_128 refers to an MX format whose element format is a custom float4 with 1 sign bit, 2-bit exponent, 1-bit mantissa, and a block size of 128.
4.3 GPTQ
We adopted an enhanced version of GPTQ compatible with MX weight formats (Sharify et al., 2024), with two additional improvements. First, we tuned the dampening factor layerwise as a hyperparameter. For each layer, we did a grid search over the space $\left\{10^{-3},10^{-2},\cdots,10^{3},10^{4}\right\}$ and chose the dampening factor that minimized layerwise output mean squared error (MSE). Second, in contrast to Frantar et al. (2022) who performed sequential layerwise Hessian accumulation and optimization to minimize GPU memory usage, we did Hessian accumulation in unquantized network for all layers before optimization. In consistency with the original work, 128 sequences from the training data split was used for Hessian accumulation.
4.4 Loss landscape mapping
All NLL losses were evaluated on the entire validation data split at half precision. Second-order loss landscape features requiring backward passes, namely Hessian-vector products, were computed in single precision using the PyHessian package (Yao et al., 2020).
5 Conclusion
Figure 6: Tradeoff between quantized model weight size and its generalization. The models in each subplot from top to bottom are: gpt2, gpt2-medium, gpt2-large, gpt2-xl; opt-125m, opt-350m, opt-1.3b, opt-2.7b, opt-6.7b, opt-13b; bloom-560m, bloom-1b1, bloom-1b7, bloom-3b, bloom-7b1; Llama-2-7b-hf, Llama-2-13b-hf, Meta-Llama-3-8B. The marker colors represent different quantized precision. Circles represent models quantized to mxfp formats, diamonds those quantized to mxint formats, with hollow markers standing for RTN and filled markers GPTQ. Black filled squares represent the pre-trained float model. Dashed/dotted gray lines connects the losses of the same model quantized to different data format families. There are 4 such lines for each model: mxint (RTN): dotted, mxfp (RTN): dotted, mxint (GPTQ): dashed, and mxfp (GPTQ): dashed. We highlight the difference before and after GPTQ by a vertical colored dashed line.
In this work, we demonstrated that, just like that of pre-training, the outcome of post-training quantization of well-trained LLMs can also be predictable, thanks to underlying scaling laws governing the local loss landscape, numerical data formats and effects of PTQ algorithms. In Figure 6, we display the tradeoff between model quantization and quality across all models and quantization formats in this study. This graph establishes a Pareto frontier, the optimal tradeoff between larger models quantized to lower bit precisions and smaller models quantized to higher bit precisions. Moreover, since our random forest model can accurately predict NLL loss after PTQ across different model sizes and distinct model families, we argue that identifying the appropriate model size and data format for a given inference workflow is no longer a business of trial-and-error but rather one of reason guided by the underlying scaling laws of quantized LLMs. Overall, we believe our findings would provide practical value to the deployment of LLMs on resource-constrained devices.
6 Limitations
Due to constraint of computational resources, we experimented with models up to 13 billion parameters. The predictive power of our scaling rules on much larger LLMs is pending further validation.
Appendix A Scaling of $\ell_{2}$-norms of model weights
In Figure 7, we summarize the scaling of the $\ell_{2}$-norms of transformer weights, for all models in the 5 LLM families under study. We found that, with the exception of the GPT-2 and OPT families, $\left\lVert{\bm{w}}\right\rVert$ scales close to half power laws w.r.t. parameter count $D$, suggesting a rather constant element-wise weight magnitude across models of different sizes. We also found that, not surprisingly, the closeness to half power law scaling of $\ell_{2}$-norms is correlated with the constancy of SQNRs for all MX data types across models.
Figure 7: Scaling of weight $\ell_{2}$-norm. Convention same as in Figure 1 Left. Light gray lines in the background mark square-root power laws, $\left\lVert{\bm{w}}\right\rVert\propto D^{\frac{1}{2}}$.
Appendix B SQNR and NLL of MX formats
Figure 8: SQNRs and NLL losses resulting from weight quantization, before PTQ. We show round-to-nearest (RTN) results for all models in multiple LLM families. Consistent with convention set in Figure 1 (left), model families are color-coded and model sizes are encoded by symbol sizes.
Figure 9: SQNRs and NLL losses resulting from weight quantization, after PTQ. Similar to Figure 8, we show GPTQ results for all models in multiple LLM families.
Appendix C Scaling in the case of PTQ to traditional int quantization
We note that, in the case of traditional weight quantization to integer (int) numerical formats, an extra step of calibration is necessary. Calibration optimizes additional parameters per quantizer, namely a scale and/or a zero point, depending on the quantization scheme. The affine transformation prescribed by the scale and zero point can also have varied granularities, from per-tensor, per-group to per-channel. Furthermore, different optimization objectives could be used to determine scale and zero point. These extra parameters and procedures likely introduce additional variability into the scaling of PTQ of LLMs, making traditional int quantization more unpredictable than MX quantization.
With concrete examples, here we show that this is indeed the case. We create and calibrate int quantizers at varied precisions and granularities, denoted by intP_(chan|gG|tens). For example, int4_tens represents a 4-bit per-tensor format, and int3_g32 a 3-bit per-group format with group size 32. We chose symmetric quantization scheme (with scale and no zero point) and calibrate by minimizing mean squared error (MSE) of quantization. Calibration data are 128 sequences taken from the training split.
Figure 10: SQNRs induced by traditional int versus MX quantizers for the smallest 3 models in the OPT family. For notations of int formats and procedural details of calibration see the main text. Numerical precision is color-coded and symbol sizes encode model capacity.
Figure 11: Scaling of SQNRs and NLL losses before and after PTQ, for int versus MX data types. Convention same as in Figure 3. Data for opt-1.3b are shown, with int and MX formats separated in 2 panels.
Not surprisingly, we find that SQNRs from int quantization are much more variable than those from MX quantization, and do not seem to scale monotonically with model size (Figure 10). In addition, the changes to SQNRs and NLL losses as a consequence of GPTQ are much less predictable in the cases of int than MX data types (Figure 11).
Appendix D Interpretation of the importance of input features to the predictive model
Beyond making accurate predictions of the difference in NLL loss between GPTQ and RTN, interpreting our predictive model can grant insight into the specific characteristics that make GPTQ most effective and the scenarios in which GPTQ should be employed.
Figure 12: Importance and interpretation of features used by our predictive model. Mean and standard deviation of the importance score (Gini importance) for each input feature, calculated across all 120 trees in the random forest (left). The predictive model’s feature-specific decision-making process for quantizing mosaicml/mpt-7b to the mxint3_64 format (right).
The Gini importance, also known as mean decrease in impurity, measures how much each feature contributes to reducing the Gini impurity in the dataset when making splits Louppe et al. [2013]. As shown in (Figure 12, left), our random forest regressor pays the most attention to the NLL loss of RTN, which can intuitively be explained by the understanding that GPTQ improves off of the baseline RTN quantization. Partial dependence graphs further reveal that the model pays more attention to the NLL loss of RTN at higher loss values, which is reasonable given that a higher starting NLL loss leaves greater room for GPTQ improvement. The number of parameters, the NLL loss of the original model, and the local loss slope are also considered by the predictive model because they describe the initial conditions of each LLM that differentiate their individual loss landscapes.
The quantization format accounts for three input features, namely precision, number of exponent bits, and block size. Of these features, precision has the largest influence on model prediction, which agrees with our findings that the largest variation in NLL loss between formats is driven by the number of bits (Figure 9, right). Note that the information gained from the quantization format is likely also embedded in the SQNR of RTN due to the strong correlation between SQNR and data format shown in (Figure 8, left), explaining why SQNR of RTN is also an important model feature.
The waterfall plot in (Figure 12, right), highlights one example of how each input feature contributes to the random forest’s prediction of the effect of GPTQ in quantizing the mosaicml/mpt-7b model to the mxint3_64 format.
Appendix E Cost of loss landscape feature computation
Figure 13: Computational cost of GPTQ versus loss landscape mapping. We show data measured from runs of 3 models from the OPT family on a single A100 GPU, where time needed for loss landscape mapping is measured on 3 random weight perturbations.
Our predictive model does not rely on features requiring second-order information, only empirical loss evaluation at critical points in the parameter space. Thus, only a few forward passes are needed to compute the input features to carry out a prediction, making the extraction of predictive features inexpensive. In Figure 13, we measure wall-clock time of feature extraction and compare it to conducting GPTQ optimization. We find that the overhead of running GPTQ is significantly more than measuring the step-wise loss landscape of 3 random weight perturbations, with the difference in overhead scaling with the model size. In practice, we only need loss landscape information local to the SNR of RTN, which could further reduce the amount of computation needed. It is much more economical to use the predictive model based on scaling, than to actually compute GPTQ.